# Robust Finetuning of Vision-Language-Action Robot Policies via Parameter Merging

### v2

@@ A.6.2 LIBERO Finetuning Datasets
A.6.3 LIBERO Evaluation Details
We use the LIBERO simulator for evaluation. Each ID eval is conducted for 20 episodes. Each OOD eval is conducted with 5 seeds, 10 episodes/seed. 20, 21, and 22 show the 3 types of OOD variations we test for in each of the 3 tasks.
@@ A.6.3 LIBERO Evaluation Details
A.7 Hyperparameters
A.7.1 Choosing Merging Coefficient $\alpha$
@@
A.7.2 Choosing Number of Gradient Steps
@@
A.7.3 Real-world Experiments’ Hyperparameters
@@
A.7.4 Simulation Experiments’ Hyperparameters
@@
A.8 Details on Baseline Methods
@@
: We fine-tune all parameters using the behavioral cloning loss, and each update batch is sampled from both the pretraining dataset and the finetuning dataset with a fixed weight. See Section˜A.7 for specific weight values we use for different tasks.
@@
A.9 Why does RETAIN work so well?
@@
A.10 Qualitative Analysis of Success and Failure Mode of RETAIN
Typically, we observe that RETAIN improves the robustness of the merged policy on OOD evaluations. Compared to task-FT policies, which are brittle and will fail in out-of-distribution scenarios catastrophically and is unable to retry, the RETAIN policies typically exhibits more robust behavior, and can recover from failure using its generalist knowledge. Typically, we observe that the task-FT policy either either does the full task successfully, or cannot do the task at all. In comparison, the RETAIN policies usually at least partially complete the task. However, we do observe that the RETAIN policies sometimes fail due to (1) imprecise execution of the task and stuck in constant retry mode and (2) produces an action that does not solve the task (though still semantically meaningful), and is unable to successfully continue afterwards. We provide two qualitative examples of this in Fig.˜23 and Fig.˜24.

### v3

@@ A.7.2 LIBERO Finetuning Datasets
A.7.3 LIBERO Evaluation Details
We use the LIBERO simulator for evaluation. Each ID eval is conducted for 20 episodes. Each OOD eval is conducted with 5 seeds, 10 episodes/seed. 21, 22, and 23 show the 3 types of OOD variations we test for in each of the 3 tasks.
@@ A.7.3 LIBERO Evaluation Details
A.8 Hyperparameters
A.8.1 Choosing Merging Coefficient $\alpha$
@@
A.8.2 Choosing Number of Gradient Steps
@@
A.8.3 Real-world Experiments’ Hyperparameters
@@
A.8.4 Simulation Experiments’ Hyperparameters
@@
A.9 Details on Baseline Methods
@@
: We fine-tune all parameters using the behavioral cloning loss, and each update batch is sampled from both the pretraining dataset and the finetuning dataset with a fixed weight. See Section˜A.8 for specific weight values we use for different tasks.
@@
A.10 Why does RETAIN work so well?
@@
A.11 Qualitative Analysis of Success and Failure Mode of RETAIN
Typically, we observe that RETAIN improves the robustness of the merged policy on OOD evaluations. Compared to task-FT policies, which are brittle and will fail in out-of-distribution scenarios catastrophically and is unable to retry, the RETAIN policies typically exhibits more robust behavior, and can recover from failure using its generalist knowledge. Typically, we observe that the task-FT policy either either does the full task successfully, or cannot do the task at all. In comparison, the RETAIN policies usually at least partially complete the task. However, we do observe that the RETAIN policies sometimes fail due to (1) imprecise execution of the task and stuck in constant retry mode and (2) produces an action that does not solve the task (though still semantically meaningful), and is unable to successfully continue afterwards. We provide two qualitative examples of this in Fig.˜24 and Fig.˜25.